# Less Is More for Multi-Step Logical Reasoning of LLM Generalisation Under Rule Removal, Paraphrasing, and Compression

### v2

@@ Abstract
Large language models (LLMs) achieve strong performance on many natural language tasks, yet their generalisation under structured perturbations of logical rule systems remains insufficiently characterised. We present a controlled evaluation framework that probes reasoning reliability through four stress tests: (1) rule deletion, removing redundant versus essential rules from a multi-step inference chain; (2) contradictory evidence injection; (3) logic-preserving rewrites based on equivalence laws (contraposition, double negation, implication-to-disjunction, De Morgan, identity, and commutativity); and (4) multi-law equivalence stacking that composes 2–5 transformations.
Across three representative model families—BERT, Qwen2, and LLaMA-like models—all models attain Acc=1.0000=1.0000 on the base split and show no degradation under redundant rule deletion. In contrast, essential rule deletion yields a pronounced decrease to near-chance performance, and injecting explicit contradictions reduces accuracy to 0.0000. Under logic-preserving rewrites, accuracy is largely preserved for single-law transformations with only small degradations in a few cases, whereas multi-law stacking exposes model-dependent sensitivity: BERT matches the base condition, TinyLlama shows only marginal degradation, and Qwen2 exhibits a substantial drop.
Overall, the results indicate that contemporary LLMs are generally stable under semantic-preserving reformulations, yet remain brittle to missing or inconsistent evidence and may degrade under composed logical transformations depending on the model family. The proposed framework provides a concise diagnostic tool for isolating these failure modes and for evaluating logical generalisation beyond surface-form variation.
Codehttps://github.com/14H034160212/lemo
Introduction
Reasoning over structured rule systems is a foundational capability for reliable decision-making and language understanding. Although large language models (LLMs) perform strongly on many benchmarks, it remains unclear to what extent such performance reflects robust logical inference as opposed to distributional heuristics. Moreover, practical reasoning settings frequently involve incomplete information, redundant premises, rephrased or logically transformed rules, and even explicit contradictions. Characterising model behaviour under these controlled perturbations is therefore important for both scientific understanding and safe deployment.
To address this gap, we introduce a controlled evaluation framework that isolates key aspects of rule-based reasoning. The benchmark is built around a canonical multi-step inference chain (e.g., green $\rightarrow$ cold $\rightarrow$ rough $\rightarrow$ young $\rightarrow$ nice) requiring multi-hop integration of rules. From each base instance, we generate targeted variants designed to probe distinct dimensions of generalisation:
Rule deletion. We remove either redundant rules (whose removal should preserve derivability) or essential rules required to complete the inference chain, thereby testing sensitivity to missing inferential links.
Contradictory evidence injection. We inject explicit facts that conflict with rule-derived conclusions, testing whether models can accommodate inconsistency rather than defaulting to a single inference path.
Logic-preserving transformations. We rewrite rules using standard equivalence laws (contraposition, double negation, implication-to-disjunction, De Morgan, identity, and commutativity). These variants preserve semantics while altering surface form, allowing us to test whether models track logical content beyond templated phrasing.
Multi-law equivalence stacking. We compose 2–5 equivalence laws simultaneously to create structurally more complex but logically equivalent rule sets, testing robustness under composed transformations.
We evaluate the framework on three model families—BERT, Qwen2, and LLaMA-like models—under identical training and evaluation pipelines. Despite architectural differences, the observed behaviour is broadly consistent:
  • Base reasoning is solved at Acc=1.0000=1.0000.
  • Redundant rule deletion yields no measurable degradation.
  • Essential rule deletion produces a marked decrease to near-chance performance.
  • Injecting explicit contradictions reduces accuracy to 0.0000.
  • Single-law equivalence rewrites are handled robustly, with only minor degradations in a small number of cases.
  • Multi-law stacking exposes model-dependent sensitivity: some models retain baseline accuracy, whereas others exhibit substantial degradation.
Together, these results highlight an asymmetry: contemporary LLMs are generally stable under semantic-preserving rewrites of rules, yet highly sensitive to missing or inconsistent evidence, and may degrade under composed logical transformations. The proposed framework provides an extensible, fine-grained diagnostic suite for isolating these failure modes, complementing existing reasoning benchmarks that often conflate multiple sources of variation.
Related Work
Research on logical reasoning in large language models (LLMs) spans multi-step deduction, logic-driven augmentation, abductive explanation, abstract reasoning assessment, generalisation analysis, and hybrid neuro-symbolic methods. A growing body of work shows that LLMs often fail even on elementary logical transformations. For example, the Reversal Curse demonstrates that models trained on “A is B” frequently fail to infer the logically equivalent “B is A,” revealing structural weaknesses in bidirectional generalisation (berglund2023reversal). Such findings motivate deeper investigation into how LLMs internalise logical relationships and under which conditions reasoning fidelity deteriorates.
Early studies like Transformers as Soft Reasoners demonstrated that transformers can perform multi-step inference over natural language rules and generalise to deeper reasoning depths (transformerClark). Subsequent work showed that this generalisation is fragile: Bao et al. found that models trained on short reasoning chains fail to extrapolate to longer or structurally shifted chains, highlighting persistent deficiencies in compositional generalisation (bao2022multi). Similar concerns are echoed in broader analyses showing that LLMs lack strong inductive biases and struggle with systematic, relational, or abstract reasoning (ijcai2024p693; Cheng2025Em; cheng2026sapp).
A parallel line of work attempts to strengthen reasoning via logic-driven augmentation. Wang et al. propose a symbolic context-extension and augmentation framework that applies equivalence laws to derive new training instances, improving model performance on ReClor and LogiQA (wang-etal-2022-logic). Bao et al. extend this idea by using Abstract Meaning Representation (AMR) to generate semantically grounded logical variants that enhance generalisation to unseen reasoning depths (bao2024). Despite their effectiveness, these methods primarily expand data distributions rather than isolating the minimal perturbations that cause reasoning failures.
Beyond deduction, abductive reasoning has been explored through frameworks such as AbductionRules, which train models to generate rules explaining surprising inputs (young-etal-2022-abductionrules). Other studies examine LLM generalisation under structured perturbations: Bao et al. show that reordering, compressing, or restructuring premises leads to substantial degradation in reasoning fidelity (baoAccess). Hybrid neuro-symbolic approaches, such as ChatLogic, integrate LLMs with Prolog-style inference mechanisms to address these limitations, although they rely heavily on external symbolic scaffolds (10650138). Complementing these evaluations, Bao’s dissertation provides a unified perspective on out-of-distribution generalisation, structural generalisation, and equivalence invariance by analysing architecture-agnostic failure modes across premise deletion, contradiction insertion, and logic-preserving rewrites (bao2025thesis).
Across these threads, prior work demonstrates meaningful progress but still lacks frameworks capable of systematically disentangling generalisation to redundant versus essential evidence, contradiction sensitivity, and invariance to different classes of logical equivalence. These unresolved limitations motivate our controlled, fine-grained evaluation framework that isolates minimal structural perturbations and reveals predictable failure modes in multi-step reasoning across LLM architectures.
Methodology
1. Synthetic Logical Environment
Base facts + rule chain generation
$\Downarrow$
2. Structural Variants
V1: Redundant rule removal
V2: Essential rule removal V3: Contradictory facts V4: Logical equivalence laws (1–5)
$\Downarrow$
3. Model Fine-Tuning & Inference
BERT / Qwen2 / TinyLlama (LoRA)
$\Downarrow$
4. Evaluation & Analysis
Accuracy, $\Delta_{\text{base}}$, variant breakdown
Figure 1: Compact system architecture of our logical reasoning generalisation framework. Each base example is transformed into structural variations (rule deletion, contradiction injection, and logic-equivalence rewrites). Models are trained on base data only and evaluated across all variants to measure generalisation and sensitivity.
The system architecture in Figure 1 outlines the full pipeline used to construct, transform, and evaluate logical reasoning tasks. The process begins with the generation of synthetic logical environments containing base facts and a multi-step rule chain designed to support deterministic deductive inferences. From each base environment, we systematically derive multiple structural variants that probe different dimensions of reasoning generalisation: (1) removal of redundant rules, (2) removal of essential rules, (3) injection of explicit contradictions, and (4) application of a wide range of logical-equivalence transformations.
These transformed instances, together with the untouched base examples, form a comprehensive evaluation suite. Models such as BERT, Qwen2, and TinyLlama are fine-tuned solely on the base cases using LoRA-based parameter-efficient adaptation. They are then tested on all structural variants to assess reasoning generalisation under controlled perturbations. The evaluation module computes accuracy and deviation from the base condition ($\Delta_{\text{base}}$), enabling fine-grained analysis of model generalisation, sensitivity to missing or conflicting information, and stability under logic-preserving rewrites.
Extended Logic Rules for Multi-Step Reasoning
This chapter explores the reasoning capabilities of AI through a comprehensive case study that demonstrates the principle of “less is more” in logical deduction. We begin with a complex base example incorporating dilemma reasoning patterns, then systematically modify it in three distinct ways to reveal different aspects of AI reasoning:
  1. 1.
    Rule Reduction: Removing redundant rules while maintaining the same conclusions
  2. 2.
    Rule Equivalence: Replace multiple rules with fewer logically equivalent rules, including the laws of contraposition, implication, double negation, De Morgan, identity, and commutativity.
  3. 3.
    Rule Interference: Adding contradictory rules that disrupt the AI’s reasoning process.
All examples use consistent predicates and individuals (Anne, green, blue, cold, rough, young) to facilitate comparison across variations. The example for each variation can be found in the appendix.
Experiment
Dataset Construction To evaluate reasoning generalisation under controlled structural perturbations, we construct a synthetic dataset consisting of multi-step logical inference chains and systematically generated variants. Each base example includes a fact specifying an entity’s attribute (e.g., “Anne is green or blue”), a rule chain of five inference steps (e.g., green → cold → rough → young → nice), four questions querying intermediate and final states, and binary ground-truth answers. From each base example, we generate eleven diagnostic variants grouped into four categories:
Rule Deletion. Variant 1: removal of a logically redundant rule. Variant 2: removal of a logically essential rule. Contradictory Evidence. Variant 3: addition of an explicit contradiction in the facts. Logic-Preserving Equivalence Laws. Six single-law variants using distinct logical equivalence transformations: contrapositive, double negation, implication, De Morgan, identity, and commutativity. Multi-Law Equivalence Stacking. A final variant consisting of 2–5 equivalence-law transformations simultaneously applied to the base rule.
The dataset contains 100 base groups, of which 80 are used for training and 20 for testing; all variants are applied to all base groups, resulting in a comprehensive evaluation suite with consistent ground truth. All transformations retain natural-language surface forms to ensure that evaluation probes reasoning rather than symbolic manipulation skills.
Models Evaluated We evaluate three representative model families commonly used in reasoning research: BERT-base (devlin2018bert): a bidirectional encoder-only architecture. Qwen2-1.5B (yang2025qwen3): a decoder-only transformer representative of modern instruction-tuned LLMs. TinyLlama-1.1B (zhang2024tinyllama): an efficient LLaMA-like decoder model trained on open datasets.
Training Procedure All models are fine-tuned using LoRA (Low-Rank Adaptation) to ensure parameter-efficient training and allow direct comparison across architectures. Training is conducted on the 80 base examples only—none of the variants are seen during training. Each question is treated as an independent binary classification instance.
Input Formatting Each training instance concatenates: [facts] + [rules] + [question] with the binary label corresponding to the ground-truth answer (“T” or “F”).
Hyperparameters To maintain comparability across models, all fine-tuning uses: Batch size: 4, Learning rate: $2\times 10^{-5}$, Epochs: 10, AdamW optimizer, Maximum sequence length: 512, and LoRA rank: 8. For decoder-only models (Qwen2 and TinyLlama), the tokenizer’s pad_token is set to the eos_token to avoid padding-related errors.
Evaluation Protocol We evaluate each model on all eleven test splits: Base test set (20 unseen base examples), Rule Deletion Variants, Contradictory Evidence Variant, Six Equivalence-Law Variants, and Multi-Law Variant.
Each test contains four questions per example, producing 80 questions for the base set and 400 questions per variant. For each input, the model predicts a binary label.
During evaluation, we additionally record: The exact facts and rules presented, the equivalence laws used (if any), Question-level predictions, and a human-readable description of the structural perturbation. This allows granular error analysis of different reasoning stressors.
We use accuracy as the primary metric, reflecting the percentage of correctly answered questions in each test split. Because each question is binary and independent, accuracy provides a direct measure of reasoning correctness.
Discussion and Implications

### v3

# Conflict-Aware Fusion: Resolving Logic Inertia in Large Language Models via Structured Cognitive Priors

@@ Abstract
Large language models (LLMs) excel at many natural language tasks, yet their reasoning reliability under structured perturbations of rule-based systems remains brittle. We present a controlled evaluation framework consisting of four stress tests: (1) rule deletion (redundant vs. essential); (2) contradictory evidence injection; (3) logic-preserving rewrites; and (4) multi-law equivalence stacking. While representative model families (BERT, Qwen2, and TinyLlama) achieve Acc=1.0000=1.0000 on base tasks, our framework reveals a critical failure mode termed Logic Inertia—a total breakdown (Acc=0.0000=0.0000) under contradictions, where deductive momentum overrides factual reality.
To resolve this, we propose Conflict-Aware Fusion, a framework grounded in the Cognitive Structure Hypothesis which posits that robust reasoning requires an explicit structural inductive bias. By imposing a dual-process architecture that separates premise verification from logical deduction, Conflict-Aware Fusion successfully eliminates logic inertia, achieving 1.0000 accuracy on both base and contradictory stress tests, and significantly enhancing robustness to missing evidence. Our results demonstrate that, for reliable multi-step reasoning, structural verification discipline is as critical as training data scale, providing a novel blueprint for building robust, contradiction-aware AI systems  [^1] https://github.com/14H034160212/lemo . See the OpenAI/Evals pull request  [^2] https://github.com/openai/evals/pull/1622 .
1 Introduction
Reasoning over structured rule systems is a foundational capability for reliable decision-making and language understanding. Although large language models (LLMs) perform strongly on many benchmarks, a fundamental question remains unresolved: do these models genuinely perform logical inference, or do they primarily rely on pattern completion learned from data distributions? This distinction is not merely theoretical. In real-world deployments—such as legal analysis, scientific discovery, and automated decision systems—reasoning agents must operate under conditions of incomplete information, redundant evidence, and explicit contradictions. A system that cannot properly validate premises before drawing conclusions risks producing confident yet logically invalid outputs. Consequently, understanding and improving the robustness of LLM reasoning under such conditions is both a scientific priority and a prerequisite for safe deployment.
However, existing evaluations of LLM reasoning largely conflate linguistic competence with logical robustness, providing limited insight into how models behave when the structural integrity of rule systems is deliberately perturbed. In particular, the field lacks diagnostic frameworks capable of isolating the effects of missing rules, redundant rules, and contradictory premises within a unified experimental setting. Without such controlled analysis, it remains difficult to determine whether current models truly reason over structured knowledge or merely emulate reasoning-like behavior.
To address this gap, we introduce a controlled evaluation framework that systematically probes the structural robustness of rule-based reasoning through four complementary stress tests: (1) Rule deletion, which distinguishes the impact of removing redundant versus essential rules; (2) Contradictory rule injection, which evaluates whether models can detect and appropriately respond to logical inconsistency; (3) Logic-preserving transformations, which assess invariance under equivalence-based reformulations of rules; and (4) Multi-law stacking, which tests compositional robustness under increasingly complex rule interactions. By isolating these factors within a unified benchmark, our framework provides a fine-grained diagnostic lens on the structural reliability of LLM reasoning.
Through a comprehensive evaluation of three representative model families—BERT, Qwen2, and TinyLlama—we uncover a previously unrecognized failure pattern that we term the Asymmetry of Robustness. Contemporary LLMs exhibit strong stability under semantics-preserving rewrites, suggesting sensitivity to surface-level logical equivalence. Yet when confronted with explicit contradictions, they display a systematic collapse in performance, yielding an accuracy of 0.0000. We characterize this phenomenon as Logic Inertia: the tendency of models to persist along learned deductive trajectories even when foundational premises become inconsistent. This finding exposes a critical blind spot in current training paradigms, which appear to optimize for completing inferential chains rather than verifying their validity. Importantly, the universality of this failure is further confirmed through external evaluation on the Human Last Exam platform [^3] https://lastexam.ai/, where all top-tier models failed our constructed cases. These results suggest that the inability to resolve logical conflict is not a marginal weakness but a structural limitation of current LLM reasoning.
Building on this diagnosis, we propose Conflict-Aware Fusion, a reasoning framework grounded in the Cognitive Structure Hypothesis. We argue that robust reasoning requires an explicit structural separation between premise validation and deductive execution—an inductive bias largely absent from current end-to-end training regimes. Our framework operationalizes this principle through a dual-process architecture in which a System 2–style contradiction detection stage precedes a System 1–style rule application stage. This design enforces structural verification before inference, thereby preventing deductive processes from proceeding over inconsistent premises. Empirical results demonstrate that this structural intervention not only resolves logic inertia but also establishes a new level of robustness across all evaluated stress conditions.
This work makes three primary contributions. First, it establishes a diagnostic benchmark that enables precise measurement of structural robustness in rule-based reasoning, disentangling the effects of redundancy, incompleteness, and contradiction. Second, it identifies and formalizes the phenomenon of Logic Inertia, revealing a fundamental limitation in current LLM reasoning that remains invisible under conventional benchmarks. Third, it introduces the Conflict-Aware Fusion framework, demonstrating that incorporating explicit premise-validation mechanisms can transform model behavior, achieving 1.0000 accuracy under both standard and contradictory conditions while substantially improving resilience to essential rule deletion (Variant 2).
More broadly, our findings challenge the prevailing assumption that scaling data and model size alone will yield reliable reasoning. Instead, they suggest that the next stage of progress in artificial reasoning systems depends on integrating structural verification mechanisms that enforce logical consistency prior to inference. By providing both a diagnostic lens and a constructive solution, this work contributes toward the development of reasoning systems that are not only powerful, but also verifiably reliable.
2 Related Work
Fragility of Logical Reasoning in LLMs
Research on logical reasoning in Large Language Models (LLMs) encompasses a broad spectrum, including multi-step deduction, abductive explanation, and abstract reasoning assessment Berglund et al. (2023); Gendron et al. (2024); Cheng et al. (2025); Xiong et al. (2025); Cheng et al. (2026). While early studies suggested that transformers could act as “soft reasoners” capable of multi-step inference Clark et al. (2021), subsequent work has revealed significant structural weaknesses. For instance, the “Reversal Curse” refers to a failure of bidirectional generalization: when a model learns facts in the form “A is B,” it often cannot infer the logically equivalent reverse fact “B is A.” For example, even though “A is the mother of B” logically implies “B is the child (or son/daughter) of A,” models trained on only one direction struggle to make this inference in the reverse direction. Berglund et al. (2023). Furthermore, Bao et al. Bao et al. (2022) demonstrate that LLMs struggle with compositional generalization, often failing to extrapolate beyond the reasoning depths or structural patterns encountered during training.
Logic-Driven Augmentation and Neuro-Symbolic Methods
To address these deficiencies, a parallel line of research explores logic-driven data augmentation and hybrid architectures. Wang et al. (2022) proposed a symbolic context-extension framework utilizing equivalence laws to augment training sets, while Bao et al. Bao et al. (2024) leveraged Abstract Meaning Representation (AMR) to generate semantically grounded logical variants. Beyond data-centric approaches, neuro-symbolic methods like ChatLogic, Wang et al. (2024) integrate LLMs with external Prolog-style engines to ensure deductive validity. However, these solutions primarily focus on expanding seen distributions or relying on external symbolic scaffolds, rather than resolving the model’s intrinsic sensitivity to minimal structural changes in the input.
Generalisation under Structured Perturbations
Recent investigations have pivoted toward analyzing model behavior under controlled perturbations, such as reordering, compressing, or restructuring premises Bao et al. (2025); Young et al. (2022); Xiong et al. (2026). These studies indicate that reasoning fidelity degrades sharply when the surface form of a problem is altered, even if the underlying logic remains invariant. Building on this, Bao’s unified perspective on out-of-distribution generalization Bao (2025) identifies architecture-agnostic failure modes across premise deletion, contradiction injection, and logical rewrites. Despite these advances, there remains a critical need for a framework capable of systematically disentangling sensitivity to redundant versus essential evidence, contradiction handling, and invariance across diverse equivalence laws. This gap motivates our fine-grained evaluation framework, which isolates minimal structural perturbations to reveal predictable failure patterns in multi-step reasoning.
3 Methodology
We introduce the Conflict-Aware Fusion framework, a structurally guided reasoning approach that enforces an explicit separation between premise verification and deductive execution. Guided by the Cognitive Structure Hypothesis, our methodology integrates three interdependent components: (1) a structural robustness benchmark for controlled evaluation, (2) a dual-process reasoning architecture that enforces verification-before-deduction, and (3) a structure-aligned optimization pipeline. Together, these components operationalize reasoning as a structurally constrained process rather than a purely data-driven behavior.
3.1 Logic Inertia as a Design Constraint
As established in Section 1, contemporary LLMs exhibit logic inertia: a tendency to persist along learned deductive trajectories even when foundational premises are incomplete or contradictory. In this work, logic inertia is treated not merely as an empirical observation but as an operational design constraint. Existing approaches that rely on contradiction exposure or large-scale data augmentation do not enforce an explicit verification stage within the reasoning trajectory, allowing deductive continuation to proceed unchecked. We therefore formalize the Cognitive Structure Hypothesis:
Reliable multi-step reasoning requires an explicit structural separation between premise verification and deductive execution.
Under this hypothesis, robust reasoning depends on the structural organization of inference rather than solely on knowledge scale or exposure. Models must first establish the logical coherence and sufficiency of premises before applying inference rules. This principle directly informs both the evaluation framework and the architecture introduced below.
3.2 Structural Robustness Benchmark
A central component of our methodology is a structural robustness benchmark designed for controlled and fine-grained evaluation of reasoning reliability. The benchmark serves a dual role: it functions as a diagnostic tool for identifying structural reasoning failures and as a structured data source for model training and optimization.
Unlike conventional benchmarks composed of heterogeneous tasks, all evaluation instances are derived from a single canonical rule-based system with systematically controlled perturbations. Each variant preserves domain semantics and vocabulary while differing only in structural properties, enabling behavioral differences to be attributed directly to reasoning robustness rather than domain shift.
Canonical reasoning backbone.
The canonical backbone defines a reference reasoning state from which all perturbed variants are derived. A concrete instantiation of this backbone is provided in the Appendix as the base example, which illustrates specific facts, rules, and deduction chains. This distinction clarifies that the benchmark is not limited to this single instance, but systematically evaluates reasoning over a unified structural template.
Controlled structural perturbations.
Benchmark splits are generated by modifying one structural property at a time while preserving semantic domain and vocabulary. Structural robustness is evaluated along four orthogonal dimensions:
  • Structural necessity sensitivity: preserve conclusions under redundant rule deletion while revising conclusions when essential rules are removed.
  • Consistency verification: detect logical contradictions prior to inference.
  • Logical-form invariance: maintain identical entailments under semantics-preserving transformations.
  • Compositional robustness: remain stable under stacked structural rewrites.
Benchmark generation (Variant 2 & 3).
We generate all instances from a single canonical rule-based template (canonical backbone) and create hard structural variants by controlled perturbations that preserve surface vocabulary and domain semantics. Given a base instance $(F,R,Q)$ (facts, rules, and a fixed set of yes/no queries), we construct: Variant 2 by removing a key rule on the main deduction chain, which breaks downstream entailments; and Variant 3 by injecting an explicit contradictory fact, which forces the model to detect inconsistency and halt/reject deduction. All splits are serialized into a unified CSV schema containing facts, rules, questions, and answers. Other variants are deferred to the Appendix.
Algorithm 1 Structural Variant Generation (Variant 2 & 3)
1:Canonical generator $\mathsf{Base}(\cdot)$, key rule $r^{\star}$, contradiction template $\mathsf{Contr}(\cdot)$, number of instances NN
2:Datasets $\mathcal{D}_{base}$, $\mathcal{D}_{v2}$, $\mathcal{D}_{v3}$
3:for $i=1$ to NN do
4:  $(F,R,Q,A)\leftarrow\mathsf{Base}(i)$ $\triangleright$ canonical backbone instance
5:  $\mathcal{D}_{base}\leftarrow\mathcal{D}_{base}\cup\{(F,R,Q,A)\}$
6:  $R_{v2}\leftarrow R\setminus\{r^{\star}\}$ $\triangleright$ Variant 2: remove an essential rule
7:  $A_{v2}\leftarrow\mathsf{Label}(F,R_{v2},Q)$
8:  $\mathcal{D}_{v2}\leftarrow\mathcal{D}_{v2}\cup\{(F,R_{v2},Q,A_{v2})\}$
9:  $F_{v3}\leftarrow F\cup\{\mathsf{Contr}(F,R)\}$ $\triangleright$ Variant 3: inject a contradiction
10:  $A_{v3}\leftarrow\mathsf{Label}(F_{v3},R,Q)$
11:  $\mathcal{D}_{v3}\leftarrow\mathcal{D}_{v3}\cup\{(F_{v3},R,Q,A_{v3})\}$
12:end for
13:Serialize $\mathcal{D}_{base},\mathcal{D}_{v2},\mathcal{D}_{v3}$ to CSV with a unified schema.
3.3 Architectural Overview: Dual-Process Reasoning
Guided by the structural failures identified above, we introduce the Conflict-Aware Fusion architecture. The core innovation is the imposition of a dual-process reasoning structure within the Chain-of-Thought (CoT) generation path.
As illustrated in Figure 1, the reasoning process is constrained such that deductive rules cannot be applied until the model has explicitly validated the logical consistency of the premises. This design converts verification from an optional behavior into a required structural step. When contradictions or missing dependencies are detected, the model halts or revises reasoning rather than continuing forward deduction.
Conflict-Aware CoT Input Facts: Socrates is a man. Rule: All men are mortal. New Fact: Socrates is NOT mortal. Instruction: ”Step 1: Check Contradiction…” SFT Logic Verify? Trace A (Correct): “Step 1: Conflict Verified! (Man $\to$ Mortal vs Not Mortal) Step 2: STOP. Trace B (Naive): “Step 1: Consistent. Step 2: Deduce Mortal.” Result: False (Halt) Result: True (Hallucinated) $\pi_{DPO}$ Verification $\pi_{SFT}$ HallucinationSystem 2 CheckSystem 1 Inertia
Figure 1: Conflict-Aware Fusion Architecture. The Prompt Template (top) imposes a structural prior requiring an explicit “Contradiction Detection” (Step 1). The model generates reasoning traces that branch based on verification. DPO (left box) reinforces the “Halt” path when contradictions are present, resolving Logic Inertia.
This architecture operationalizes the Cognitive Structure Hypothesis by embedding a verification-first inductive bias directly into the reasoning process, ensuring that inference is based on validated premises.
3.4 Structure-Aligned Optimization Pipeline
To ensure consistent adherence to the proposed structural prior, we implement a two-stage optimization pipeline that explicitly reinforces verification-before-deduction as a learned reasoning constraint.
Stage 1: Structural SFT.
We first perform supervised fine-tuning on a balanced dataset of 11,200 instances spanning canonical, perturbed, and contradiction-containing variants. All training samples enforce a mandatory “Step 1: Verify facts” preamble, requiring the model to explicitly assess premise completeness and consistency before attempting deduction. By uniformly applying this structure across both base and robustness-focused variants, the training process discourages shortcut learning on familiar patterns and promotes generalizable verification behavior. This stage establishes premise checking as a default procedural step rather than an optional response to rare anomalies.
Stage 2: Logical Alignment via DPO.
We further refine the model using Direct Preference Optimization (DPO) to align generation behavior with structural reliability. Preference pairs are constructed to contrast verification-compliant reasoning with opportunistic deductive continuation. Responses that correctly trigger a “Halt Reasoning” state upon detecting contradictions or insufficient premises are consistently preferred over those that proceed with unsupported inference. This alignment step directly penalizes hallucinated shortcuts and reinforces disciplined termination when logical validity cannot be established.
Through this combined pipeline, premise verification is internalized as a governing structural constraint that shapes how reasoning unfolds rather than merely influencing final answers. The model learns not only to produce correct conclusions but to regulate the conditions under which deduction is permitted, thereby reducing logic inertia and improving robustness under structural perturbations.
4 Experiment and Results
4.1 Main Results: Conflict-Aware Fusion Performance
The primary objective of our evaluation is to assess whether the proposed framework preserves logical integrity under controlled structural stressors. As reported in Table 1, all baseline models indicate strong performance under unperturbed conditions. However, this apparent reliability collapses once the structural integrity of the rule system is disrupted. When essential rules are removed (Variant 2), accuracy drops sharply across all models, and under explicit contradictions (Variant 3), performance uniformly falls to 0.0000. These results provide direct evidence of pronounced logic inertia.